If you were to titrate 49.6 mL of 0.036 M solution of HCl with a 0.046 M solution of NaOH, what volume of NaOH in mL would be required to reach the equivalence point?
1
Expert's answer
2020-11-23T07:39:50-0500

The equivalence point (in titrimetric analysis) — point of titration when the number of equivalents of titrant added is equivalent, or equal to the number of equivalents of the analyte in the sample.

"С1\timesМ1=C2\times V2"

"0.036\times49.6=0.046\times V2"

V2=38.8